1.   A square metal plate with a length of 0.20 m and with a charge of 9.7 x 10 -7 C. Find the electric field of the square metal plate. 


1
Expert's answer
2021-04-19T09:05:09-0400

Answer

Electric field due to square plate

"E=\frac{\sigma }{2\epsilon _0}\\=\frac{q}{2A\epsilon _0}\\=\frac{9.7\times10^{-7}}{2\times0.2\times0.2\times8.85 \times10^{-12} }\\=1.4\times10^{6}N/C"